# Mingchen Li1, Anne Blaes2, Steven Johnson3, Hongfang Liu4, Hua Xu5, Rui Zhang1

### v1

@@ Abstract
Medical Large Language Models (LLMs) such as ClinicalCamel 70B, Llama3-OpenBioLLM 70B have demonstrated impressive performance on a wide variety of medical NLP task. However, there still lacks a large language model (LLM) specifically designed for cancer domain. Moreover, these LLMs typically have billions of parameters, making them computationally expensive for healthcare systems. Thus, in this study, we propose CancerLLM, a model with 7 billion parameters and a Mistral-style architecture, pre-trained on 2,676,642 clinical notes and 515,524 pathology reports covering 17 cancer types, followed by fine-tuning on three cancer-relevant tasks, including cancer phenotypes extraction, cancer diagnosis generation, and cancer treatment plan generation. Our evaluation demonstrated that CancerLLM achieves state-of-the-art results compared to other existing LLMs, with an average F1 score improvement of 8.1%. Additionally, CancerLLM outperforms other models on two proposed robustness testbeds. This illustrates that CancerLLM can be effectively applied to clinical AI systems, enhancing clinical research and healthcare delivery in the field of cancer.
@@ 1 Introduction
Figure 1: The evolution of medical LLM performance on three tasks—cancer phenotype extraction, diagnosis generation, and treatment plan generation—is measured using the average F1 score, which includes Exact Match, BLEU-2, and ROUGE-L. Our CancerLLM achieves the highest performance with an F1 score of 78.00%.
Bridging these gaps, in this work, we introduce CancerLLM, a state-of-the-art LLM with 7 billion parameters designed to elevate LLM’s proficiency in the cancer domain, and aid medical professionals in cancer phenotype extraction, cancer diagnosis generation, and cancer treatment plan generation. Our model represents an important step towards democratizing advancements in the cancer domain of LLMs. More specifically, our research focuses on four steps: 1) pre-training the LLM in the style of Mistral. We pre-train the Mistral 7B using 2,676,642 cancer clinical notes and 515,524 pathology reports from the University of Minnesota (UMN) Clinical Data Repository. 2) fine-tuning the LLM on a specialized dataset. We proposed three datasets focusing on phenotype extraction, diagnosis generation and treatment plan generation. Following this, we fine-tuned CancerLLM using instruction learning for the three aforementioned tasks. 3) We evaluate the generation ability of CancerLLM against other medical LLMs using Exact Match (Rajpurkar et al., 2016), BLEU-2 (Papineni et al., 2002), and ROUGE-L (Cohan & Goharian, 2016) metrics. 4) To further evaluate the model’s ability, we propose two robustness testbeds which include counterfactual robustness and misspellings robustness in the special tasks. As shown in Figure  1, our model not only outperforms existing 7B and 13B models by a significant margin but also delivers good results with those of 70B models.
@@
  • An original study with the introduction of a benchmark of three tasks, facilitating assessment against existing state-of-the-art open-source medical LLMs.
  • We have developed three datasets aimed at fine-tuning CancerLLM for its three downstream tasks.
@@ 2.1 Pre-training Data for Cancer
The data used in this study were obtained from the University of Minnesota (UMN) Clinical Data Repository. It contains the health records of 31,465 cancer patients. Specifically, It includes 2,676,642 cancer clinical notes (7.27GB) and 515,524 cancer pathology reports (536.85MB). We obtained the data with the approval of the UMN Institutional Review Board (IRB). The statistics information of patient and clinical notes are shown in Figure 3.
(a) Number of patients with different cancers
(b) Number of clinical notes with different cancers
Figure 3: Patients and clinical notes Statistics
@@ 2.2.2 Diagnosis Generation
Dataset We randomly selected 10,635 cancer clinical notes that do not overlap with the pre-training dataset, and split them into 80% instances for training, 10% instances for testing, and 10% instances for validation. These datasets do not overlap. Each instance follows the template {instruction, context, response}. The instruction is: You are a medical expert. This task involves generating the diagnosis based on the provided context or text, the context includes the reason for visit, treatment site, subjective information, nursing Review of Systems (ROS), objective observations, and laboratory test results. The response is the cancer diagnosis.
2.2.3 Treatment Plan Generation
Task description By giving the information from clinical notes, which includes the 1) reason for visit, 2) treatment site, 3) subjective information, 4) nursing Review of Systems (ROS), 5) objective observations, 6) laboratory test results, 7) diagnosis and 8) assessment, the model is expected to generate the correct treatment plan.
Dataset We randomly selected 10,635 clinical notes that do not overlap with the pre-training dataset, and split them into 80% instances for training, 10% instances for testing, and 10% instances for validation. These datasets do not overlap. Each instance follows the template {instruction, context, response}. The instruction is: You are a medical expert. This task involves generating the treatment plan based on the provided context or text, the context includes the reason for visit, treatment site, subjective information, nursing Review of Systems (ROS), objective observations, laboratory test results, diagnosis, and assessment. The response is the treatment plan.
@@ 3.1.1 Cancer Diagnosis Generation
To assess the scalability of our model, in this part, we evaluate the performance of CancerLLM in diagnosis generation task. Table 2 presents the experiment results of various medical LLMs. We have the following observations: (1) our CancerLLM significantly outperforms all the strong baselines and its variants across all evaluation metrics. (2) We observed that CancerLLM improve the original Mistral 1*7B, and Bio-Mistral 7B by 28.93%, and 17.92% respectively, in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (3) Our model not only achieves the best performance among 7B models but also surpasses the performance of medical LLMs with 13B and 70B parameters. (4) Without domain knowledge, larger parameter LLMs are ineffective. For instance, Llama3-OpenBioLLM-70B and ClinicalCamel-70B achieve only 62.81% and 63.55%, separately.
3.1.2 Treatment Plan Generation
@@
PMC LLaMA 7B 10.61 10.61 10.61 64.52 64.52 64.52 74.08 74.08 74.08 49.74
Medalpaca 7B 11.11 1.11 11.11 65.26 65.26 65.26 74.69 74.69 74.69 50.35
LLAMA-2 7B 8.51 8.51 8.51 62.28 62.28 62.28 72.09 72.09 72.09 47.63
Mistral 1*7B 14.01 14.01 14.01 67.30 67.30 67.30 75.94 75.94 75.94 52.42
Mixtral 8*7B 12.91 12.91 12.91 65.83 65.83 65.83 74.98 74.98 74.98 51.24
Bio-Mistral 7B 13.71 13.71 13.71 67.85 67.85 67.85 76.61 76.61 76.61 52.72
LLama3 8B 12.31 12.31 12.31 65.60 65.60 65.60 74.26 74.26 74.26 50.72
MedLLaMA 13B 11.91 11.91 11.91 66.48 66.48 66.48 75.59 75.59 75.59 51.33
PMC LLaMA 13B 12.11 12.11 12.11 66.34 66.34 66.34 75.62 75.62 75.62 51.13
Medalpaca 13B 12.11 12.11 12.11 65.66 65.66 65.66 74.55 74.55 74.55 50.77
LLaMA2 13B 9.51 9.51 9.51 64.36 64.36 64.36 73.93 73.93 73.93 49.27
LLaMA2 70B 14.01 14.01 14.01 66.55 66.55 66.55 75.61 75.61 75.61 52.06
Llama3-OpenBioLLM-70B 9.01 9.01 9.01 62.76 62.76 62.76 72.54 72.54 72.54 48.10
ClinicalCamel-70B 14.01 14.01 14.01 67.23 67.23 67.23 75.97 75.97 75.97 52.40
CancerLLM(Ours) 15.72 15.72 15.72 71.09 71.09 71.09 79.41 79.41 79.41 55.40
Table 3: Results of Cancer Treatment Plan Generation
Table 3 presents the experiment results of various approaches on the treatment plan generation task. We have the following observations: (1) our CancerLLM significantly outperforms all the strong baselines and its variants across all evaluation metrics. (2) We observed that CancerLLM improve the original Mistral 1*7B, and Bio-Mistral 7B by 2.59%, and 2.68% respectively, in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (4) Without domain knowledge, larger parameter LLMs still are ineffective. For instance, Llama3-OpenBioLLM-70B and ClinicalCamel-70B achieve only 48.10% and 52.40%, separately. (5) An LLM with more parameters is not always better than one with fewer parameters. For instance, MedLLaMA 13B achieves a performance of 51.33%, while Llama3-OpenBioLLM-70B only reaches 48.10%. (6) Current medical LLMs struggle to generate an exact plan that matches the one outlined in clinical notes created by doctors, as evidenced by the exact match results in Table 3. Despite our model achieving the best performance, it only attains a 15.72% Exact Match (F1) score.
@@ 3.2.2 Misspellings Robustness
Diagnosis Generation Bio-Mistral 7B 2% 0.82 0.82 0.82 10.38 10.38 10.38 25.23 25.23 25.23 10.18
@@
20% 0.09 0.09 0.09 11.09 11.09 11.09 27.09 27.09 27.09
40% 0.00 0.00 0.00 10.13 10.13 10.13 27.16 27.16 27.16
60% 0.00 0.00 0.00 3.75 3.75 3.75 18.42 18.42 18.42
80% 0.00 0.00 0.00 1.16 1.16 1.16 16.25 16.25 16.25
Treatment Plan Generation Bio-Mistral 7B 2% 0.00 0.00 0.00 22.77 22.77 22.77 37.00 37.00 37.00 14.03
4% 0.00 0.00 0.00 17.64 17.64 17.64 31.61 31.61 31.61
6% 0.00 0.00 0.00 18.09 18.09 18.09 32.25 32.25 32.25
8% 0.00 0.00 0.00 16.60 16.60 16.60 30.27 30.27 30.27
20% 0.00 0.00 0.00 16.62 16.62 16.62 30.29 30.29 30.29
40% 0.00 0.00 0.00 14.59 14.59 14.59 27.76 27.76 27.76
60% 0.00 0.00 0.00 14.28 14.28 14.28 27.17 27.17 27.17
CancerLLM(Ours) 2% 0.00 0.00 0.00 23.97 23.97 23.97 38.16 38.16 38.16 15.58
4% 0.00 0.00 0.00 18.00 18.00 18.00 31.92 31.92 31.92
6% 0.00 0.00 0.00 19.87 19.87 19.87 34.24 34.24 34.24
8% 0.00 0.00 0.00 20.92 20.92 20.92 35.07 35.07 35.07
20% 0.00 0.00 0.00 17.34 17.34 17.34 30.47 30.47 30.47
40% 0.00 0.00 0.00 18.80 18.80 18.80 32.15 32.15 32.15
60% 0.00 0.00 0.00 19.84 19.84 19.84 33.28 33.28 33.28
Table 6: Misspellings robustness performance on diagnosis generation and treatment plan generation. The rate refers to different misspelling rates in each instance.
To validate the robustness of our model, we propose a misspelling robustness testbed designed to simulate misspelling in the clinical notes. Table 6 presents the complete results and average F1 performance of different counterfactual rates across exact match, BLEU-2, and ROUGE-L metrics for both diagnosis generation and treatment generation tasks. Our observations are as follows: (1) CancerLLM consistently outperforms Bio-Mistral 7B. For the diagnosis generation task, CancerLLM achieves an average F1 score of 10.66%, compared to Bio-Mistral 7B’s 10.18%. In the treatment plan generation task, CancerLLM scores 15.58%, while Bio-Mistral 7B scores 14.03%. For both diagnosis generation and treatment plan generation tasks, when the counterfactual rate is set to 2%-80%, the performance is nearly zero.
@@ 3.3 Results of Generation Time and GPU Memory
Phenotype Extraction Treatment Plan Generation Diagnosis Generation
LLMs F1 Time Used Memory F1 Time Used Memory F1 Time Used Memory
Bio-Mistral 7B 91.24 1:06:55 5,746 MB 52.72 9:04:56 5,786 MB 68.89 1:07:45 5,802 MB
Mistral 1*7B 91.65 57:05 5,598 MB 52.42 8:57:08 5,650MB 55.07 1:07:49 5,680 MB
Mixtral 8*7B 92.26 2:01:27 25,086 MB 51.24 11:57:42 2,5128 MB 60.12 2:16:14 25,166 MB
PMC LLaMA 13B 90.06 1:08:59 8,208 MB 51.13 12:09:34 10,012 MB 62,86 1:19:52 9,208 MB
LLaMA2 13B 91.86 1:08:43 8,204 MB 49.27 12:08:25 10,090 MB 55.42 1:24:17 9,254 MB
ClinicalCamel-70B 93.72 2:50:16 37,716 MB 52.40 25:58:21 37,670 MB 63.55 3:05:37 37,67 MB
CancerLLM(Ours) 91.78 1:14:12 5,550 MB 55.40 11:40:14 5738 MB 86.81 1:26:33 5,768 MB
@@
In Table 7, we compared the generation time on the whole testing set and used GPU memory on a single A100 across different LLMs. For the Phenotype Extraction task, we set the maximum input token length to 1500 and the maximum new token length to 50. For the Treatment Plan Generation task, we set the maximum input token length to 4000 and the maximum new token length to 500. In the Diagnosis Generation task, we set the maximum input token length to 1500 and the maximum new token length to 500. We observed that although Mixtral 8*7B, LLaMA2 13B, and ClinicalCamel-70B exhibit better phenotype extraction performance, they have higher generation time and GPU memory usage. For example, ClinicalCamel-70B requires 2:50:16 for inference and uses 37,716 MB of GPU memory, while CancerLLM requires only 1:14:12 for inference and uses just 5,550 MB of GPU memory.
@@ 4.1 Cancer Diagnosis Generation
In this task, we primarily pre-train CancerLLM and explore its effectiveness in the diagnosis generation task by providing relevant information. As shown in Table 2, the Bio-Mistrial 7B obtains the best performance among all baseline LLMs. We guess that one reason for this is that Bio-Mistral 7B is trained on a large pre-processed corpus, such as the PubMed Central corpus Canese & Weis (2013). Another reason is that the foundational structure of Mixtral enhances training effectiveness. The Mixtral 8*7B outperforms the Mixtral 1*7B by more than 5% in average F1 value. However, Mixtral 8*7B is a Mixture of Experts (MoE) model with 8 experts, making training and inference time challenging. Fortunately, our CancerLLM achieves the best model performance through pre-training and fine-tuning with a single expert model. Not all LLMs with large parameters achieve the best performance. For example, the 70B models (LLaMA2 70B, Llama3-OpenBioLLM-70B, and ClinicalCamel-70B) do not outperform the Bio-Mistral 7B, despite being trained on medical corpora. This indicates that smaller parameter LLMs can achieve better performance. Our 7B model, CancerLLM, which incorporates cancer domain knowledge, also supports this point. This is particularly important in the medical domain as it reduces training time and memory usage, providing opportunities for medical institutions with limited computational resources.
4.2 Cancer Treatment Plan Generation
By comparing extract match results of Table 3 and Table 2, we found that LLMs struggle with treatment plan generation, as even the best model, CancerLLM, achieves only 15.72% performance. One reason is that the length of the ground truth treatment plan is longer than the diagnosis, making it difficult to generate the same treatment plan as created by doctors. Mistral 1*7B is also the most powerful model among all baselines, outperforming even the 70B models such as ClinicalCamel-70B. Mistral 1*7B achieves 52.42% average F1 performance, whereas ClinicalCamel-70B achieves only 50.52% average F1. It also provides evidence that smaller LLMs could achieve better performance compared to larger models. Increasing model parameters is not the sole method for enhancing model performance; exploring how to incorporate high-quality domain knowledge is also crucial.

### v2

# Mingchen Li1, Jiatan Huang1, Jeremy Yeung1, Anne Blaes2 Steven Johnson3, Hongfang Liu4, Hua Xu5, Rui Zhang1

@@ Abstract
Medical Large Language Models (LLMs) such as ClinicalCamel 70B, Llama3-OpenBioLLM 70B have demonstrated impressive performance on a wide variety of medical NLP task. However, there still lacks a large language model (LLM) specifically designed for cancer domain. Moreover, these LLMs typically have billions of parameters, making them computationally expensive for healthcare systems. Thus, in this study, we propose CancerLLM, a model with 7 billion parameters and a Mistral-style architecture, pre-trained on 2,676,642 clinical notes and 515,524 pathology reports covering 17 cancer types, followed by fine-tuning on two cancer-relevant tasks, including cancer phenotypes extraction, and cancer diagnosis generation. Our evaluation demonstrated that CancerLLM achieves state-of-the-art results compared to other existing LLMs, with an average F1 score improvement of 7.61%. Additionally, CancerLLM outperforms other models on two proposed robustness testbeds. This illustrates that CancerLLM can be effectively applied to clinical AI systems, enhancing clinical research and healthcare delivery in the field of cancer.
@@ 1 Introduction
Figure 1: The evolution of medical LLM performance on cancer phenotype extraction, and diagnosis generation is measured using the average F1 score, which includes Exact Match, BLEU-2, and ROUGE-L. Our CancerLLM achieves the highest performance with an F1 score of 74.54%.
Bridging these gaps, in this work, we introduce CancerLLM, a state-of-the-art LLM with 7 billion parameters designed to elevate LLM’s proficiency in the cancer domain, and aid medical professionals in cancer phenotype extraction, cancer diagnosis generation. Our model represents an important step towards democratizing advancements in the cancer domain of LLMs. More specifically, our research focuses on four steps: 1) pre-training the LLM in the style of Mistral. We pre-train the Mistral 7B using 2,676,642 cancer clinical notes and 515,524 pathology reports from the University of Minnesota (UMN) Clinical Data Repository. 2) fine-tuning the LLM on a specialized dataset. We proposed three datasets focusing on phenotype extraction, and diagnosis generation. Following this, we fine-tuned CancerLLM using instruction learning for the aforementioned tasks. 3) We evaluate the generation ability of CancerLLM against other medical LLMs using Exact Match (Rajpurkar et al., 2016), BLEU-2 (Papineni et al., 2002), and ROUGE-L (Cohan & Goharian, 2016) metrics. 4) To further evaluate the model’s ability, we propose two robustness testbeds which include counterfactual robustness and misspellings robustness in the special tasks. As shown in Figure  1, our model not only outperforms existing 7B and 13B models by a significant margin but also delivers good results with those of 70B models.
@@
  • An original study with the introduction of a benchmark of phenotype extraction and diagnosis generation, facilitating assessment against existing state-of-the-art open-source medical LLMs.
  • We have developed three datasets aimed at fine-tuning CancerLLM for its two downstream tasks.
@@ 2.1 Pre-training Data for Cancer
The data used in this study were obtained from the University of Minnesota (UMN) Clinical Data Repository. It contains the health records of 31,465 cancer patients. Specifically, It includes 2,676,642 cancer clinical notes (7.27GB) and 515,524 cancer pathology reports (536.85MB). We obtained the data with the approval of the UMN Institutional Review Board (IRB).
@@ 2.2.2 Diagnosis Generation
Dataset We randomly selected 10,635 cancer clinical notes that do not overlap with the pre-training dataset, and split them into 80% instances for training, 10% instances for testing, and 10% instances for validation. These datasets do not overlap. Each instance follows the template {instruction, context, response}. The instruction is: You are a medical expert. This task involves generating the diagnosis based on the provided context or text, the context includes the reason for visit, treatment site, subjective information, nursing Review of Systems (ROS), objective observations, and laboratory test results. The response is the cancer diagnosis. In general, when recording a diagnosis in the clinical notes, doctors may describe the same diagnosis using different terms or abbreviations, such as ’L lung cancer’ to refer to ’left lung cancer.’ To standardize these diagnoses, ICD codes are introduced. However, mapping these varied descriptions to a unified ICD code is a challenging task. To evaluate our model’s performance, we also created a test dataset containing 374 clinical notes, where the ground truth diagnosis is represented by the ICD code stored in the database. This testing dataset is called ICDdiagnosis.
@@ 3.1.1 Cancer Diagnosis Generation
To assess the scalability of our model, in this part, we evaluate the performance of CancerLLM in diagnosis generation task. Table 2 presents the experiment results of various medical LLMs. We have the following observations: (1) our CancerLLM significantly outperforms all the strong baselines and its variants across all evaluation metrics. (2) We observed that CancerLLM improve the original Mistral 1*7B, and Bio-Mistral 7B by 28.93%, and 17.92% respectively, in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (3) Our model not only achieves the best performance among 7B models but also surpasses the performance of medical LLMs with 13B and 70B parameters. (4) Without domain knowledge, larger parameter LLMs are ineffective. For instance, Llama3-OpenBioLLM-70B and ClinicalCamel-70B achieve only 62.81% and 63.55%, separately. (5) Table 3 presents the experiment results of various medical LLMs on ICDdiagnosis, We observed that CancerLLM improve the best best baseline model Bio-Mistral 7B by 4.36% in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (6) The 70B models do not achieve the best performance.
@@
PMC LLaMA 7B 24.33 24.33 24.33 35.66 35.66 35.66 44.66 44.66 44.66 34.90
Medalpaca 7B 27.27 27.27 27.27 38.09 38.09 38.09 47.86 47.86 47.86 37.73
LLAMA-2 7B 24.06 24.06 24.06 35.02 35.02 35.02 44.36 44.36 44.36 34.48
Mistral 1*7B 26.74 26.74 26.74 39.00 39.00 39.00 48.53 48.53 48.53 38.09
Mixtral 8*7B 29.68 29.68 29.68 38.18 38.18 38.18 47.86 47.86 47.86 38.57
Bio-Mistral 7B 30.21 30.21 30.21 41.16 41.16 41.16 50.69 50.69 50.69 40.68
LLama3 8B 27.01 27.01 27.01 38.16 38.16 38.16 48.05 48.05 48.05 37.73
MedLLaMA 13B 29.41 29.41 29.41 39.80 39.80 39.80 48.34 48.34 48.34 39.18
PMC LLaMA 13B 26.74 26.74 26.74 37.02 37.02 37.02 46.03 46.03 46.03 36.59
Medalpaca 13B 25.13 25.13 25.13 35.57 35.57 35.57 45.31 45.31 45.31 35.33
LLaMA2 13B 22.99 22.99 22.99 34.08 34.08 34.08 43.42 43.42 43.42 33.49
LLaMA2 70B 27.81 27.81 27.81 38.85 38.85 38.85 47.94 47.94 47.94 38.20
Llama3-OpenBioLLM-70B 27.01 27.01 27.01 39.39 39.39 39.39 49.35 49.35 49.35 38.58
ClinicalCamel-70B 28.34 28.34 28.34 38.75 38.75 38.75 48.05 48.05 48.05 38.37
CancerLLM(Ours) 34.49 34.49 34.49 46.03 46.03 46.03 54.62 54.62 54.62 45.04
Table 3: Results of Cancer Diagnosis Generation on ICDdiagnosis
@@ 3.2.2 Misspellings Robustness
Diagnosis Generation Bio-Mistral 7B 2% 0.82 0.82 0.82 10.38 10.38 10.38 25.23 25.23 25.23 12.06
@@
ICDdiagnosis Bio-Mistral 7B 2% 0.00 0.00 0.00 7.07 7.07 7.07 18.00 18.00 18.00 8.62
4% 0.00 0.00 0.00 7.27 7.27 7.27 17.78 17.78 17.78
6% 0.27 0.27 0.27 8.07 8.07 8.07 18.63 18.63 18.63
8% 0.00 0.00 0.00 7.65 7.65 7.65 18.79 18.79 18.79
CancerLLM(Ours) 2% 0 0 0 8.93 8.93 8.93 20.59 20.59 20.59 9.57
4% 0.27 0.27 0.27 9.07 9.07 9.07 20.53 20.53 20.53
6% 0.27 0.27 0.27 8.34 8.34 8.34 19.34 19.34 19.34
8% 0 0 0 8.20 8.20 8.20 19.31 19.31 19.31
Table 6: Misspellings robustness performance on diagnosis generation and ICDdiagnosis. The rate refers to different misspelling rates in each instance.
To validate the robustness of our model, we propose a misspelling robustness testbed designed to simulate misspelling in the clinical notes. Table 6 presents the complete results and average F1 performance of different counterfactual rates across exact match, BLEU-2, and ROUGE-L metrics for both diagnosis generation and ICDdiagnosis. Our observations are as follows: CancerLLM consistently outperforms Bio-Mistral 7B. For the diagnosis generation task, CancerLLM achieves a similar average F1 score compared to Bio-Mistral 7B. Specifically, in ICD diagnosis, CancerLLM scores 9.57%, while Bio-Mistral 7B scores 8.62%.
@@ 3.3 Results of Generation Time and GPU Memory
Phenotype Extraction Diagnosis Generation1
LLMs F1 Time Used Memory F1 Time Used Memory
Bio-Mistral 7B 91.24 1:06:55 5,746 MB 68.89 1:07:45 5,802 MB
Mistral 1*7B 91.65 57:05 5,598 MB 55.07 1:07:49 5,680 MB
Mixtral 8*7B 92.26 2:01:27 25,086 MB 60.12 2:16:14 25,166 MB
PMC LLaMA 13B 90.06 1:08:59 8,208 MB 62,86 1:19:52 9,208 MB
LLaMA2 13B 91.86 1:08:43 8,204 MB 55.42 1:24:17 9,254 MB
ClinicalCamel-70B 93.72 2:50:16 37,716 MB 63.55 3:05:37 37,67 MB
CancerLLM(Ours) 91.78 1:14:12 5,550 MB 86.81 1:26:33 5,768 MB
@@
In Table 7, we compared the generation time on the whole testing set and used GPU memory on a single A100 across different LLMs. For the Phenotype Extraction task, we set the maximum input token length to 1500 and the maximum new token length to 50. In the Diagnosis Generation task, we set the maximum input token length to 1500 and the maximum new token length to 500. We observed that although Mixtral 8*7B, LLaMA2 13B, and ClinicalCamel-70B exhibit better phenotype extraction performance, they have higher generation time and GPU memory usage. For example, ClinicalCamel-70B requires 2:50:16 for inference and uses 37,716 MB of GPU memory, while CancerLLM requires only 1:14:12 for inference and uses just 5,550 MB of GPU memory.
@@ 4.1 Cancer Diagnosis Generation
In this task, we primarily pre-train CancerLLM and explore its effectiveness in the diagnosis generation task by providing relevant information. As shown in Table 2, the Bio-Mistrial 7B obtains the best performance among all baseline LLMs. We guess that one reason for this is that Bio-Mistral 7B is trained on a large pre-processed corpus, such as the PubMed Central corpus Canese & Weis (2013). Another reason is that the foundational structure of Mixtral enhances training effectiveness. The Mixtral 8*7B outperforms the Mixtral 1*7B by more than 5% in average F1 value. However, Mixtral 8*7B is a Mixture of Experts (MoE) model with 8 experts, making training and inference time challenging. Fortunately, our CancerLLM achieves the best model performance through pre-training and fine-tuning with a single expert model. Not all LLMs with large parameters achieve the best performance. For example, the 70B models (LLaMA2 70B, Llama3-OpenBioLLM-70B, and ClinicalCamel-70B) do not outperform the Bio-Mistral 7B, despite being trained on medical corpora. This indicates that smaller parameter LLMs can achieve better performance. For ICDdiagnosis, we find that overall performance is lower than in diagnosis generation. This is because the expected diagnoses in ICD diagnosis are more complex. Our 7B model, CancerLLM, which incorporates cancer domain knowledge, also supports this point. This is particularly important in the medical domain as it reduces training time and memory usage, providing opportunities for medical institutions with limited computational resources.